• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name fansler/ewing

    Log entry time 15:34:37 on September 11, 2009

    Entry number 289134

    keyword=Cryo/Power Supply Faults

    The left Q1 liquid level meter has been replaced. The liquid level was behaving erratically. We noticed the lead flows were low so R. Norton adjusted the lead flow capacity and all lead flow have been adjusted to spec. There were multiple magnet trips and resets as a result of the initial LQ1 trip. All power supplies have been reset.